OpenStudy (anonymous):

Write an expression in simplest form for the volume of a cube with a side of 11 x to the 6th power a.121x^12 b.1331x^18 c.121x^36 d.1331x^12

OpenStudy (anonymous):

\((11x^6)^3=11^3\times (x^6)^3\)

OpenStudy (anonymous):

\(11^3=1331\) and \((x^6)^3=x^{6\times 3}=x^{18}\)

OpenStudy (anonymous):

you got that?